# PDP: Physics-Based Character Animation via Diffusion Policy

### v1

@@ IV Experimental Details
We demonstrate the generality and effectiveness of PDP by applying it to three distinctive applications using three different datasets: locomotion control under large physical perturbations using addbiomechanics dataset [7], universal motion tracking using AMASS [8], and physics-based text-to-motion synthesis using the KIT subset of AMASS and HumanML3D [9] for text labels. Experimental details for each application are described in this section, with results, comparisons, and ablation studies presented in the next section.
IV-A Perturbation Recovery
@@
The goal of the perturbation recovery task is to train a single diffusion policy that is capable of capturing the wide range of human responses to perturbations. Being able to model and simulate this behavior is important for studying human robustness [28] and for designing better exo-skeleton or prosthetic systems [29].
IV-A1 Dataset
@@
We use the Bump-em dataset, a subset of the Addbiomechanics dataset [7], which consists of recovery motions of human participants being physically pushed while walking on a treadmill. Participants are perturbed in the same stance with varying forces and directions applied to the hip through a parallel tethered robot [30]. The recorded motions demonstrate that participants exhibit a diverse set of recovery strategies even under the same perturbation and initial stance, which makes this dataset particularly well suited for studying how well a model can capture the multi-modality of human behaviour.
@@
For this task, trials from one participant were used. Perturbations were collected as the subject walked forward on a treadmill at a fixed speed, impacted at left toe-off stance in four directions (front, left, right, back) with magnitudes of $7.5\%$ and $15\%$ body weight. Two trials of each configuration were collected, resulting in 16 motions for the participant. However, two were dropped due to poor motion quality resulting in a dataset of 14 motion trajectories.
IV-A2 Experimental Details
@@
We use a 25-joint skeletal model from Addbiomechanics [7], optimized for the specific participant’s inertia and joint lengths. The environment is simulated with Mujoco and consists of the simulated treadmill and the skeletal model. In this environment, the observation space is defined in the world frame. The RL agent’s observation consists of the bodies center of mass positions $\bm{x}^{p}\in\mathbb{R}^{3B}$ and linear velocities $\dot{\bm{x}}^{p}\in\mathbb{R}^{3B}$, as well as the bodies rotation ${\bm{x}}^{r}\in\mathbb{R}^{3\times 3\times B}$ During RL training, the agent receives the same perturbation experienced by the human during the trial and optimizes tracking the human response through the collected reference motion following [10].
@@
After training expert RL policies for each motion, we collect new observations for PDP, by including a binary signal $\bm{p}$ that indicates if the human is being perturbed, without detailing the force magnitude or direction. This signal is helpful for the diffusion policy to differentiate between normal walking and perturbation recovery; otherwise, the policy would react to non-existent perturbations. This addition is justified, as humans can discern disturbances directly, providing no predictive advantage. The full perturbation recovery observation can then be defined as a tuple $(\bm{x}^{p},\bm{x}^{r},\dot{\bm{x}}^{p},\dot{\bm{x}}^{r},\bm{p})$. 14 motions were sampled using the various noisy/clean state action strategies, and used to train PDP. Each model was trained on a single RTX 2080 Ti GPU for approximately 1.5 hours.
TABLE I: Performance with different sampling strategies
@@
State Action Success (%) $\uparrow$ $E_{\text{g-mpjpe}}$downarrow $E_{\text{mpjpe}}$downarrow $E_{\text{vel}}$downarrow $E_{\text{acc}}$downarrow Success ID (%) $\uparrow$ FPV $\downarrow$
@@
IV-B Universal Motion Tracking
@@
TABLE II: Motion tracking results on AMASS train and test datasets
@@
IV-B1 Dataset
We use the AMASS dataset, and use the same train/test splits as PHC [14], where the subsets “Transitions Mocap” and “SSM synced” are used for testing and the rest are placed in the training set. Since our method is agnostic to the RL policies themselves, we utilize pre-trained motion tracking controller PHC [14] to track most of the motions and train individual RL policies for challenging motions where PHC fails [10]. With PHC and a few specialized small policies, our expert policies produce a training set that covers most motions in AMASS [8] and KIT [31]. Following the same practice in PHC, we exclude motions that are infeasible in our physics simulators, such as leaning on tables.
IV-B2 Experimental Details
We use the humanoid model from [14], which follows the SMPL kinematic structure with $J=23$ spherical joints. The reference motion includes the linear position $\bm{x}^{p}_{ref}\in\mathbb{R}^{3J}$ and 6d rotation $\bm{x}^{r}_{ref}\in\mathbb{R}^{6J}$ of each joint. The linear and angular velocities are found by finite difference $\dot{\bm{x}}^{p}_{ref}\in\mathbb{R}^{3J}$, $\dot{\bm{x}}^{r}_{ref}\in\mathbb{R}^{3J}$ respectively. The full motion tracking observation can then be defined as a tuple $(\Delta\bm{x}^{p},\Delta\bm{x}^{r},\Delta\dot{\bm{x}}^{p},\Delta\dot{\bm{x}}^{r},\bm{x}^{p}_{ref},\bm{x}^{r}_{ref})$, where $\Delta\bm{x}^{p}=\bm{x}^{p}_{ref}-\bm{x}^{p}$ and the other $\Delta$ terms are similarly defined. All quantities are measured in the character frame, where the origin is placed at the root of the character, the x-axis aligns with the character’s facing direction, and the z-axis points upward.
@@ IV-C Text-to-Motion
IV-C1 Dataset
For training data, we use KIT dataset and the text annotations from HumanML3D [9]. The task vector is generated by passing the text annotation through the CLIP embedding [26]. We use the same pre-trained tracking controller, PHC, to obtain the observations and actions for training.
IV-C2 Experimental Details
@@ V Results
Figure 3: PDP rollouts for a 15% body weight perturbation where the white pelvis and arrows indicate where the force is applied and the direction, respectively. Each row demonstrates a unique mode of recovery from the same perturbation.
(a) Ground truth distribution
(b) VAE $\beta$ 0.01
(c) VAE $\beta$ 0.001
(d) PDP horizon 1
(e) PDP horizon 6
Figure 4: Global left foot contact positions after $15\%$ body weight perturbation in meters. +Y and +X align with the character’s forward and right directions, respectively. The different colored arrows represent the directions that the force is applied on the person. The shaded areas represent foot contacts in the training distribution with noise level 0.12. The black X’s represent the ground truth foot contacts of the human participant. All policies were trained on the stochastic dataset with noise level 0.12.
V-A Sampling Strategy
We examine the impact of sampling strategies on the tracking and perturbation tasks. For the tracking task, training was conducted using the KIT dataset, with evaluations performed on the test set employed in the AMASS-Test split. Table I records the quantitative results of different sampling strategies.
@@ V-A Sampling Strategy
Noisy-state clean-action strategy outperforms other strategies, achieving a perfect success rate of $100\%$ in the perturbation task and $93.5\%$ success rate in the tracking task, see Table I. These results underscore the importance of selecting the right sampling strategy. By visiting out-of-distribution states and collecting the optimal action, this approach generates higher quality demonstrations, leading to better generalization and robustness.
V-B Perturbation Recovery
@@
We compare the performance of PDP to two other baselines, a C-VAE and an MLP. The selection of these approaches was based on their reliance on supervised learning principles. The C-VAE introduces an alternative generative model framework, whereas the MLP serves as a deterministic alternative. The C-VAE uses a similar setup as [18] where state and next state are fed into an encoder to produce a latent code. A decoder takes a randomly sampled latent vector alongside the current state to produce the current action. Both the C-VAE and MLP follow the same architecture as PDP with minor algorithm-dependent adjustments such as excluding the diffusion timestep embedding.
V-B1 Robustness
@@
Robustness is measured by the successful completion of an episode, defined as the agent not falling within 6 seconds of the perturbation, a period that allows for several gait cycles to complete post-perturbation. Perturbations are categorized into In-Distribution (ID), which are the same as those in the training data, and Out-of-Distribution (OOD) perturbations, which determines the policy’s capacity to effectively handle unforeseen perturbations by adjusting aspects such as timing, intensity, and direction of impact. We sample OOD perturbations as follows: We choose a random perturbation to cover all gait phases by randomly choosing an impact timing within [0, 2] seconds (equivalent to  2.5 gait cycles overlap), a random force magnitude between $7.5\%$ and $15\%$ of body weight which represents the extrema of the forces used in the Bump-em dataset, and a random force direction that is parallel to the ground.
@@
Table III show the ID and OOD performance of each baseline. All three models can handle ID perturbations, achieving a success rate of $100\%$. However, when faced with OOD perturbations, C-VAE and PDP methods exhibit notably higher performance, with C-VAE achieving a success rate of $91.3\%$ and PDP achieving a success rate of $96.3\%$. Handling OOD impacts is challenging because the model may not have seen them before, especially the impact timing, as all training examples occur at the left toe-off. The OOD distribution performance results could be attributed to the multi-modal nature of the dataset, where using an MLP would result in policies that return the average of the response recorded in the dataset, causing the policy to fail, while C-VAE and PDP can synthesize a more tailored response from the multimodal distribution.
V-B2 Foot Placement Variability
@@
Foot placement holds significant importance in understanding perturbation response and balance [32, 33]. A reliable model should ideally mirror the human response or closely approximate it while also accounting for multi-modality. Figure 4a illustrates both the actual foot placements of the participant and the distribution of foot placements obtained through a noisy sampling procedure. When examining foot placement variability for impacts within the distribution, a lower variance in foot placement should be anticipated. This is gauged by assessing the variance in the L2 distance between each foot placement and the nearest ground truth data point. We design a metric based on this called Foot Placement Variability (FPV) to answer how spread apart the foot positions are from the policy compared to the closest ground truth position:
$\text{FPV}=\frac{1}{N}\sum_{i=1}^{N}\left(\min_{j\in\{1,2,\ldots,M\}}\sqrt{(x_{i}-\bar{x}_{j})^{2}+(y_{i}-\bar{y}_{j})^{2}}\right)$ (3)
@@
TABLE III: Baseline comparisons for Perturbation Task
Method Beta Value Success ID (%) $\uparrow$ Success OOD (%) $\uparrow$ FPV $\downarrow$
@@
TABLE IV: Noise Level and Horizon Ablation for Perturbation Task
Noise Level Horizon Success ID (%) $\uparrow$ Success OOD (%) $\uparrow$ FPV $\downarrow$
@@
V-C Motion Tracking
We demonstrate that PDP is capable of reliably tracking a significant portion of the motion in AMASS. Our method achieves a $96.4\%$ success rate on the AMASS test dataset, where failure is defined as in PHC [14]: an episode is considered a failure if at any point during evaluation the joints are, on average, more than 0.5 meters from the reference motion. In addition to the success rate, we also adopt metrics used by [14], specifically mean per-joint position error ($E_{\text{mpjpe}}$) and global mean per-joint position error ($E_{\text{g-mpjpe}}$), which assess the model’s accuracy in matching the reference motion in local and global frames. We also measure the error in velocity ($E_{\text{vel}}$) and acceleration ($E_{\text{acc}}$) between the simulated character and the motion capture data. As shown in Table II, PDP matches or outperforms PHC in all metrics.
@@ V-C Motion Tracking
V-D Text-to-Motion
Figure 1 (middle) shows our results in the text-to-motion domain. We demonstrate that PDP is capable of following diverse text commands in natural language, such as jumping and kicking commands. Evaluating a model that employs auto-regressive inference during simulation with a limited history for composite actions presents significant challenges. Specifically, a prompt such as “walk then jump” cannot be effectively executed because the agent lacks the necessary memory of the initial action. To address this, we evaluate the model using a set of 42 action text prompts from the dataset, such as “a person dances” and “a person walks forward”. Each prompt is considered successful if the model performs the specified action without falling.
@@ VI Discussion
Given the recent robotics community’s interest in developing robust humanoid locomotion policies, our findings are particularly relevant [34, 35, 36]. While a single optimal strategy might suffice for a specific impact or perturbation, our results indicate that an MLP, which cannot capture different modes, lacks robustness to out-of-distribution (OOD) perturbations. In contrast, our diffusion model effectively stores a variety of strategies, providing it with a broader base of information to draw from when an OOD impact occurs. This capability could enhance locomotion policies’ ability to handle diverse and unpredictable real-world perturbations.
@@
Previous literature has shown the difficulty of producing a single and reliable motion tracker [37, 14]. Our method can train a model directly through supervised learning and exceed the performance of more complex hierarchical RL policies. Furthermore, This capability facilitates the creation of a pre-trained tracking controller that can be swiftly adapted to new datasets. This feature is notably distinct from conventional RL-based methods [13, 14] which necessitates retraining or fine-tuning controllers. Our method requires only a few noisy rollouts and a few epochs of additional learning to obtain new skills. Given the small data volume, this process is relatively quick, taking only a few minutes excluding the training of the expert policy.
@@
Despite its advantages, our approach has notable limitations. The primary constraint is the speed of the denoising process. It is significantly slower than a single forward pass of a typical MLP, making it challenging for real-time motion tracking applications. However, recent works such as [38, 4] offer promising avenues for speeding up this process.
VII Conclusion and Future Work
We present a novel framework for physics-based character control that leverages the capability of diffusion models to capture diverse behaviors. Our proposed sampling strategy, noisy-state-clean-action, significantly outperforms alternative sampling strategies. For the perturbation recovery task, our method effectively captures the distribution of human responses and demonstrates robustness to both in-distribution and out-of-distribution perturbations. In universal motion tracking, our method surpasses the state-of-the-art performance, including our baseline using a non standard non-generative model. Additionally, we showcase our methods ability to synthesize motion conditioned on text. Future work could look into speeding up the inference by using methods like [38, 39], or leveraging the large pre-trained motion tracker for downstream RL tasks.

### v2

@@ 4. Experiments
We demonstrate the generality and effectiveness of PDP by applying it to three distinctive applications using three different datasets: locomotion control under large physical perturbations using addbiomechanics dataset (Werling et al., 2023), universal motion tracking using AMASS (Mahmood et al., 2019), and physics-based text-to-motion synthesis using the KIT subset of AMASS and HumanML3D (Guo et al., 2022) for text labels. Experimental details for each application are described in this section, with results, comparisons, and ablation studies presented in the next section. Note that although each application uses the same model architecture in Figure 2 (except for Block A and B), they are trained separately with different distilled datasets.
4.1. Perturbation Recovery
@@
The goal of the perturbation recovery task is to train a single diffusion policy that is capable of capturing the wide range of human responses to perturbations. Being able to model and simulate this behavior is important for studying human robustness (Jensen et al., 2023) and for designing better exo-skeleton or prosthetic systems (Hodossy and Farina, 2023).
4.1.1. Dataset
@@
We use the Bump’em dataset, a subset of the Addbiomechanics dataset (Werling et al., 2023), which consists of recovery motions of human participants being physically pushed while walking on a treadmill. Participants are perturbed in the same stance with varying forces and directions applied to the hip through a parallel tethered robot (Tan et al., 2020). The recorded motions demonstrate that participants exhibit a diverse set of recovery strategies even under the same perturbation and initial stance, which makes this dataset particularly well suited for studying how well a model can capture the multi-modality of human behaviour.
@@
For this task, trials from one participant were used. Perturbations were collected as the subject walked forward on a treadmill at a fixed speed, impacted at left toe-off stance in four directions (front, left, right, back) with magnitudes of $7.5\%$ and $15\%$ body weight, resulting in 16 total motions; two were dropped due to poor motion quality.
4.1.2. Experimental Details
@@
We use a 25-joint skeletal model from Addbiomechanics (Werling et al., 2023), optimized for the specific participant’s inertia and joint lengths. The environment is simulated with Mujoco and consists of the simulated treadmill and the skeletal model. In this environment, the observation space is defined in the world frame. The RL agent’s observation consists of the bodies center of mass positions $\bm{x}^{p}\in\mathbb{R}^{3B}$ and linear velocities $\dot{\bm{x}}^{p}\in\mathbb{R}^{3B}$, as well as the bodies rotation ${\bm{x}}^{r}\in\mathbb{R}^{3\times 3\times B}$ During RL training, the agent receives the same perturbation experienced by the human during the trial and optimizes tracking the human response through the collected reference motion following (Peng et al., 2018).
@@
After training expert RL policies for each motion, we collect new observations for PDP, by including a binary signal $\bm{p}$ that indicates if the human is being perturbed, without detailing the force magnitude or direction. This signal is helpful for the diffusion policy to differentiate between normal walking and perturbation recovery; otherwise, the policy would react to non-existent perturbations. This addition is justified, as humans can discern disturbances directly. Furthermore, this inclusion provides no predictive advantage, similar to providing foot contact information. The full perturbation recovery observation can then be defined as a tuple $(\bm{x}^{p},\bm{x}^{r},\dot{\bm{x}}^{p},\dot{\bm{x}}^{r},\bm{p})$. 15 motions were sampled using the various noisy/clean state action strategies, and used to train PDP. Each model was trained on a single RTX 2080 Ti GPU for approximately 1.5 hours.
Table 1. Performance with different sampling strategies. The Tracking task was trained on the KIT subset of the AMASS train set and evaluated on the AMASS test set. The Perturbation task was trained on the Bump’em dataset. *Indicates our method.
@@
State Action Success (%) $\uparrow$ $E_{\text{g-mpjpe}}$downarrow $E_{\text{mpjpe}}$downarrow $E_{\text{vel}}$downarrow $E_{\text{acc}}$downarrow Success ID (%) $\uparrow$ FPC $\downarrow$
@@
4.2. Universal Motion Tracking
@@
Table 2. Motion tracking results on AMASS train and test datasets.
@@
4.2.1. Dataset
We use the AMASS dataset, and use the same train/test splits as PHC (Luo et al., 2023), where the sub-datasets: Transitions Mocap and SSM synced are used for testing and the rest are placed in the training set. Since our method is agnostic to the RL policies themselves, we utilize pre-trained motion tracking controller PHC (Luo et al., 2023) to track most of the motions and train individual policies for challenging motions where PHC fails (Peng et al., 2018). With PHC and a few specialized small policies, our training set covers most motions in AMASS (Mahmood et al., 2019) and KIT (Plappert et al., 2016). Following the same practice in PHC, we exclude motions that are infeasible in our physics simulators, such as leaning on tables.
4.2.2. Experimental Details
We use the humanoid model from (Luo et al., 2023), which follows the SMPL kinematic structure with $J=23$ spherical joints. The reference motion includes the linear position $\bm{x}^{p}_{ref}\in\mathbb{R}^{3J}$ and 6d rotation $\bm{x}^{r}_{ref}\in\mathbb{R}^{6J}$ of each joint. The linear and angular velocities are found by finite difference $\dot{\bm{x}}^{p}_{ref}\in\mathbb{R}^{3J}$, $\dot{\bm{x}}^{r}_{ref}\in\mathbb{R}^{3J}$ respectively. The full motion tracking observation can then be defined as a tuple $(\Delta\bm{x}^{p},\Delta\bm{x}^{r},\Delta\dot{\bm{x}}^{p},\Delta\dot{\bm{x}}^{r},\bm{x}^{p}_{ref},\bm{x}^{r}_{ref})$, where $\Delta\bm{x}^{p}=\bm{x}^{p}_{ref}-\bm{x}^{p}$ and the other $\Delta$ terms are similarly defined. All quantities are measured in the character frame, where the origin is placed at the root of the character, the x-axis aligns with the character’s facing direction, and the z-axis points upward.
@@ 4.3. Text-to-Motion
4.3.1. Dataset
For training data, we use KIT dataset and the text annotations from HumanML3D (Guo et al., 2022). The task vector is generated by passing the text annotation through the CLIP embedding (Radford et al., 2021). We use the same pre-trained tracking controller, PHC, to obtain the observations and actions for training.
4.3.2. Experimental Details
@@ 5. Results
5.1. Sampling Strategy
We examine the impact of sampling strategies on the tracking and perturbation tasks. For the tracking task, training was conducted using the KIT dataset, with evaluations performed on the test set employed in the AMASS-Test split. Table 1 records the quantitative results of different sampling strategies.
@@ 5.1. Sampling Strategy
Noisy-state clean-action strategy outperforms other strategies, achieving a perfect success rate of $100\%$ in the perturbation task and $93.5\%$ success rate in the tracking task, see Table 1. These results underscore the importance of selecting the right sampling strategy. By visiting out-of-distribution states and collecting the optimal action, this approach generates higher quality demonstrations, leading to better generalization and robustness.
5.2. Perturbation Recovery
@@
We compare the performance of PDP to two other baselines, a C-VAE and an MLP. The selection of these approaches was based on their reliance on supervised learning principles. The C-VAE introduces an alternative generative model framework, whereas the MLP serves as a deterministic alternative. The C-VAE uses a similar setup as (Won et al., 2022) where state and next state are fed into an encoder to produce a latent code. A decoder takes a randomly sampled latent vector alongside the current state to produce the current action. Both the C-VAE and MLP follow the same architecture as PDP with minor algorithm-dependent adjustments such as excluding the diffusion timestep embedding.
5.2.1. Robustness
@@
Robustness is measured by the successful completion of an episode, defined as the agent not falling within 6 seconds of the perturbation, a period that allows for several gait cycles to complete post-perturbation. Perturbations are categorized into In-Distribution (ID), which are the same as those in the training data, and Out-of-Distribution (OOD) perturbations, which determines the policy’s capacity to effectively handle unforeseen perturbations by adjusting aspects such as timing, intensity, and direction of impact. We sample OOD perturbations as follows: We choose a random perturbation to cover all gait phases by randomly choosing an impact timing within [0, 2] seconds (equivalent to  2.5 gait cycles overlap), a random force magnitude between $7.5\%$ and $15\%$ of body weight which represents the extrema of the forces used in the Bump’em dataset, and a random force direction that is parallel to the ground.
@@
Table 3 shows the ID and OOD performance of each baseline. All three models can handle ID perturbations, achieving a success rate of $100\%$. However, when faced with OOD perturbations, C-VAE and PDP methods exhibit notably higher performance, with C-VAE achieving a success rate of $91.3\%$ and PDP achieving a success rate of $96.3\%$. Handling OOD impacts is challenging because the model may not have seen them before, especially the impact timing, as all training examples occur at the left toe-off. The OOD distribution performance results could be attributed to the multi-modal nature of the dataset, where using an MLP would result in policies that return the average of the response recorded in the dataset, causing the policy to fail, while C-VAE and PDP can synthesize a more tailored response from the multimodal distribution.
Two important hyper-parameters in our method are the choice of noise level for creating the stochastic dataset and the action prediction horizon. Note that the action prediction horizon refers to the number of actions being predicted during training. For inference, all policies re-plan after every step of the simulation. Table 4 shows the performance across different choices of noise level and horizon. Importantly, we see that 0 noise (equivalent to clean-state clean-action) performs extremely poorly with an ID success rate of just 3.36%. Adding noise increases robustness significantly before eventually slightly harming performance. For the action prediction horizon, we note that lower horizons yield better performance, with horizon 1 achieving the best ID and OOD success rates of 100.0% and 96.3%, respectively. Larger horizons see a dramatic drop in performance. We speculate that this is because actions closer to the current timestep are more important for stability, but our loss function does not weight this importance.
5.2.2. Foot Placement Correctness
@@
Foot placement holds significant importance in understanding perturbation response and balance (Perry and Srinivasan, 2017; Rebula et al., 2013). A reliable model should ideally mirror the human response or closely approximate it while also accounting for multi-modality. Figure 4a illustrates both the actual foot placements of the participant and the distribution of foot placements obtained through a noisy sampling procedure. When examining foot placement correctness for impacts within the distribution, a lower variance in foot placement should be anticipated. This is gauged by assessing the variance in the L2 distance between each foot placement and the nearest ground truth data point. We design a metric based on this called foot placement correctness (FPC) to answer how spread apart the foot positions are from the policy compared to the closest ground truth position:
(3) $\text{FPC}=\frac{1}{N}\sum_{i=1}^{N}\left(\min_{j\in\{1,2,\ldots,M\}}\sqrt{(x_{i}-\bar{x}_{j})^{2}+(y_{i}-\bar{y}_{j})^{2}}\right)$
@@
Table 3. Baseline comparisons on the Bump-em Dataset. ID stands for in-distribution, OOD stands for out-of-distribution, and FPC stands for foot placement correctness.
Method Beta Value Success ID (%) $\uparrow$ Success OOD (%) $\uparrow$ FPC $\downarrow$
@@
Table 4. Noise Level and Horizon Ablation for Bump’em Perturbation Task. FPC is measured on the left foot contact positions.
Noise Level Horizon Success ID (%) $\uparrow$ Success OOD (%) $\uparrow$ FPC $\downarrow$
@@
5.3. Motion Tracking
We demonstrate that PDP is capable of reliably tracking a significant portion of the motion in AMASS. Our method achieves a $96.4\%$ success rate on the AMASS* test dataset, where failure is defined as in PHC (Luo et al., 2023): an episode is considered a failure if at any point during evaluation the joints are, on average, more than 0.5 meters from the reference motion. In addition to the success rate, we also adopt metrics used by (Luo et al., 2023), specifically mean per-joint position error ($E_{\text{mpjpe}}$) and global mean per-joint position error ($E_{\text{g-mpjpe}}$), which assess the model’s accuracy in matching the reference motion in local and global frames. We also measure the error in velocity ($E_{\text{vel}}$) and acceleration ($E_{\text{acc}}$) between the simulated character and the motion capture data. As shown in Table 2, PDP matches or outperforms PHC in all metrics.
@@ 5.3. Motion Tracking
5.4. Text-to-Motion
Figure 1 (middle) shows our results in the text-to-motion domain. We demonstrate that PDP is capable of following diverse text commands in natural language, such as jumping and kicking commands. Evaluating a model that employs auto-regressive inference during simulation with a limited history for composite actions presents significant challenges. Specifically, a prompt such as ”walk then jump” cannot be effectively executed because the agent lacks the necessary memory of the initial action. To address this, we evaluate the model using a set of 42 action text prompts from the dataset, such as ”a person dances” and ”a person walks forward.” Each prompt is considered successful if the model performs the specified action without falling.
@@ 6. Discussion
Given the recent robotics community’s interest in developing robust humanoid locomotion policies, our findings are particularly relevant. (Kaymak et al., 2023; Li et al., 2021; Singh et al., 2023). While a single optimal strategy might suffice for a specific impact or perturbation, our results indicate that an MLP, which cannot capture different modes, lacks robustness to out-of-distribution (OOD) perturbations. In contrast, our diffusion model effectively stores a variety of strategies, providing it with a broader base of information to draw from when an OOD impact occurs. This capability could enhance locomotion policies’ ability to handle diverse and unpredictable real-world perturbations.
@@
Previous literature has shown the difficulty of producing a single and reliable motion tracker (Won et al., 2020b; Luo et al., 2023). Our method can train a model directly through supervised learning and exceed the performance of more complex hierarchical RL policies. Furthermore, This capability facilitates the creation of a pre-trained tracking controller that can be swiftly adapted to new datasets. This feature is notably distinct from conventional RL-based methods (Won et al., 2020b, a; Luo et al., 2023) which necessitates finetuning low-level controllers or training completely new ones. Additionally, as hierarchical systems accumulate new low-level controllers, the composer faces increasing complexity, often necessitating a complete system retrain to reduce the number of low-level controllers. In contrast, our method only requires standard finetuning procedures to integrate a new dataset after training the local experts, enabling more efficient scalability with additional motion datasets.
@@
Despite its advantages, our approach has notable limitations. The primary constraint is the speed of the denoising process. Compared to the inference time of the MLP baseline, the diffusion model takes $K$ times longer, where $K$ is the number of denoising steps. This can make it challenging for applications that require high frequency control. Recent methods can reduce the number of inference steps required (Yin et al., 2023; Huang et al., 2024).
Another limitation lies in the trade-off inherent in diffusion-based policies when predicting over longer horizons. Although capturing multi-modality necessitates considering future actions, focusing on predicting multiple actions and weighting them equally can dilute the importance of the immediate action, thereby reducing robustness. This tension between long-horizon prediction for diversity and the accuracy of immediate actions is a crucial challenge. Future work could explore adaptive weighting schemes during training that balance effective long-horizon prediction with the precision of immediate actions.
7. Conclusion and Future Work
We present a novel framework for physics-based character control that leverages the capability of diffusion models to capture diverse behaviors. Our proposed sampling strategy, noisy-state-clean-action, significantly outperforms alternative sampling strategies. For the perturbation recovery task, our method effectively captures the distribution of human responses and demonstrates robustness to both in-distribution and out-of-distribution perturbations. In universal motion tracking, our method surpasses the state-of-the-art performance, including our baseline using a non standard non-generative model. Additionally, we showcase our methods ability to synthesize motion conditioned on text. Future work could look into speeding up the inference by using methods like (Yin et al., 2023; Gu and Dao, 2023), or leveraging the large pre-trained motion tracker for downstream RL tasks.